Find the exponential function of the given form that contains the given points.
Form: f(x)= c*(a^x)
Points: (0,4) and (4, 324)
I know the answer is f(x)=4*3^x,
Please explain, thanks!!

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



I already answered this for you. Stop clogging up the system with double postings.

John

My calculator said it, I believe it, that settles it

Answer by stanbon(75887)

Find the exponential function of the given form that contains the given points.
Form: f(x)= c*(a^x)
Points: (0,4) and (4, 324)
I know the answer is f(x)=4*3^x,
Form: y = c*a^x
----------------------
Using (4,324) and (0,4) you get:
324 = c*a^4
4 = c*a^0
-------
Note:: c = 4
-----
Substitute and solve for "c":
324 = 4*a^4
81 = a^4
a = 3
------
Therefore y = 4*3^x
------------------------
Cheers,
Stan H.
----------------